Elaine has two suitcases that are
similar rectangular solids. The
length of the second suitcase is
twice the length of the first
suitcase. If the first suitcase has
a volume of 1568 cubic inches,
what is the volume of the
second suitcase in cubic inches?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Elaine has two suitcases that are
similar rectangular solids. The
length of the second suitcase is
twice the length of the first
suitcase. If the first suitcase has
a volume of 1568 cubic inches,
what is the volume of the
second suitcase in cubic inches?
--------------
1st suitcase: lwh = 1568
---
2nd suitcase:(2l)(wh) 2(1568) = 3136 cu in
============================================
Cheers,
Stan H.

Answer by richwmiller(17219)

if the two solids are similar all measurements are proportional.
so if the length is double so is the width and height
2^3=8 so the volume is 8 times the first not just double.
2*L*2W*2H=8*L*W*H
8*1568=volume of second sutcase